OpenStudy (anonymous):

Quadrilateral OPQR is inscribed inside a circle as shown below. What equation would be needed to solve for angle P? What is the measure of angle P?

OpenStudy (anonymous):

Any help is very much appreciated, I'd like for someone to explain to me, not just give me an answer because I don't understand this. Thank you.

OpenStudy (anonymous):

I got 43, is this correct?

zepdrix (zepdrix):

The interior angles of a `quadrilateral` add up to 360 degrees. So we can add up our angle measures and set it equal to 360.\[\Large\rm (2x)^o+y^o+(3y+8)^o+(2x+4)^o=360^o\] But if we're not given a value for x, I'm not sure how we can get a numerical solution for y. Hmmm, what steps did you take to get 43?

OpenStudy (anonymous):

These were my steps: Angle P + angle R = 180 degrees. Angle P = y and angle R = 3y + 8. y + 3y + 8 = 180 4y + 8 = 180 angle p = 43 @zepdrix

zepdrix (zepdrix):

Opposite angles of a quadrilateral inscribed in a circle are supplementary? Ooo interesting! I didn't know that :) Your steps look correct! Yay good job!

OpenStudy (anonymous):

Thank you :3